OpenStudy (anonymous):

Is the line through the points R(-1,3) and S(2,-7) parallel to the graph of the line given by the equation, 10x+3y=6? Explain why. A.Yes, both lines are vertical B.Yes, both lines have the same slope. C.No, both lines have positive slopes that are not equal. D.No, both lines have negative slopes that are not equal.

OpenStudy (danjs):

you want to compare the slopes of both of those lines, the equation, and the line through those 2 points

OpenStudy (danjs):

so for the equation 10x + 3y = 6 The slope is: 3y = -10x + 6 y = -10/3 + 2 slope = -10/3

OpenStudy (danjs):

for the points: slope = (-7 - 3) / (2-(-1)) = -10/3

OpenStudy (danjs):

same slopes